the sum of two integers is fifty seven. Twice the smaller integer is six less than the larger integer. Find the smaller integer
Answer by hamsanash1981@gmail.com(151)

Let the Smaller and larger integers be x and y
y = 2x+6
x+Y = 57
Substituting the value of y in the above equation,
x+ 2x+6 = 57
3x =57-6
3x=51
x = 17
the smaller integer is 17